how much water must be evaporated from 240 pounds of a solution that is 3% alcohol to strengthen it to a solution that is 5% alcohol?

Answer by Alan3354(69443)

how much water must be evaporated from 240 pounds of a solution that is 3% alcohol to strengthen it to a solution that is 5% alcohol?
--------------
3% of 240 pounds = 7.2 pounds of alcohol
If 7.2 pounds = 5%, the total = 7.2/0.05 = 144 pounds
240 - 144 = 96 pounds of water to evaporate.
---------------
PS That won't work, the alcohol has a lower boiling so it will evaporate first.
If it did work, a lot of people would make stronger booze that way.